. A permanent magnet is made of ferromagnetic material with a magnetization equals to 7.5 10 A/ m 5  . The magnet is in the form of a cylindrical road with a length of 6.00 cm and a diameter of 1.20 cm. Determine the magnetic dipole moment of the magnet.


1
Expert's answer
2022-01-23T16:16:21-0500

"m=IS=\frac{\pi}4\dot I d^2 h=5.1~A\cdot m^2."